Fermi National Laboratory
 
Operation Summary for January 18, 2012 - January 20, 2012
 

Wednesday January 18

The day shift began with beam off. Operations had the Linac, Booster, Main Injector (MI), NuMI, and MiniBooNE areas ready for access.

During the day shift, at 8:31 AM, a MI person reported a LCW leak coming from a quadrupole magnet (E:Q861); Water Group personnel began repairs at 9:46 AM. Linac came out of access at 10:17 AM. Interlock personnel reported at 1:10 PM that they had completed a test of the Meson safety system. Linac began running study pulses at 3:02 PM. The Transfer Hall was out of access by 3:10 PM. NuMI was out of access by 3:13 PM. The Switchyard (SY) and Meson line were ready for beam by 3:34 PM.

During the evening shift, Operators reported a possible vacuum leak in the MI-8 line; Vacuum personnel investigated, but found no problems. MI came out of access at 4:45 PM. At 5:23 PM, Operators received a report of water flowing out of a bus duct at MI-10; Water Group people investigated. Booster came out of access at 5:36 PM. Operators reported a vacuum burst in P1 transfer line and later one in the MI-8 line; everything recovered. Operators prepared MI for access at 6:41 PM to further investigated the water leak at the MI-10 service building. Operations established beam to the Booster dump at 7:16 PM. Water Group personnel reported that they had tightened a valve and came out of MI at 7:44 PM. Water Group person reported at 7:56 PM that the Tevatron sector C1 LCW pump was leaking and turned the pump off. Operations established beam to NuMI at 8:29 PM and to MiniBooNE at 8:45 PM. At 9:06 PM, Operations reported the NuMI beam permit was tripping off due to bad beam positions; the MI Operations Specialist came in to help with tuning. Booster personnel completed their tuning at 10:51 PM; they reported that a Booster RF station (BRF1) was off and would stay off for the night.

Thursday January 19

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 1:03 AM, the Neutrino fixed target beam line NS2 LCW system tripped off on a low expansion tank indication: Operators added water to the system, but noticed oil in the basin under the pump. When they turned the pump on it made a lot of noise; they turned off the pump and later in the morning reported it to the Water Group. At 1:34 AM, the NuMI horn power supply tripped off; it reset. At 2:50 AM, a Linac Klystron RF station (KRF1) tripped off; it reset. Operations completed a tune of the H- Source and the 750 KeV line at 3:23 AM. At 4:33 AM, the NuMI beam permit tripped off due to a RF gap incomplete indication; everything reset. Operators arch and secured MTA, at 5:31 AM; this was done in preparation for beam later in the day. At 7:54 AM, Neutron Therapy Facility personnel began preparing to treat two patients later in the day.

During the day shift, at 8:25 AM, the NuMI beam permit began tripping off due to the horn power supply tripping off; this was the backup supply not the main power supply, but it was tied to the safety system. Power supply personnel investigated. Operations loaded protons into the Recycler at 9:29 AM. Operations resumed sending beam to NuMI at 10:13 AM. Interlock personnel reported at 10:25 AM that they had completed a test of the MTest (MT6) area safety system. A Water Group person turned on the NS2 LCW system at 1:43 PM. Operations began sending beam to MTA for studies at 2 PM. At 3:09 PM, Operations turned off Booster and began preparing the area for access to repair BRF3. Operations reported at 3:04 PM that SY personnel had been taking study beam during the afternoon in preparation for the next FTBF Meson experiment. Booster came out of access at 3:54 PM.

During the evening shift, Operations resumed sending beam to NuMI and MiniBooNE at 4:09 PM. Operations disabled the MTA beam at 5:15 PM. Operations began sending beam to the new Meson FTBF experiment T-1018 at 6:37 PM. Operations disabled T-1018 beam at 10:01 PM. Operators reported that they had completed making some small Booster tune changes at 10:12 PM. The FESS Duty person reported at 10:15 PM that the MI-60 North power supply room air handler had tripped off; Operators got permission for him to access the room and lock off the handler. At 10:20 PM, the MI North anode power supply tripped off.

Friday January 20

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, Operators reported at 4:18 AM that they had completed a Booster injection tune. Operators reported at 6:59 AM that they had completed a tune of the H- Source and the 750 KeV line. At 7:39 AM, NTF personnel began preparing to treat two patients later in the morning and early afternoon.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The OpsCo agrees with the plans by Booster personnel to work on their efficiencies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/20/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory